# Datalog-Expressibility for Monadic and Guarded Second-Order Logic

### v2

@@ 6.6 Separation in MSO
Since $\operatorname{dom}(h)\leq k$ it is impossible that $r=0$ and $s=k-1$; hence, if $x_{0}$ and $x_{k}$ are from $\operatorname{dom}(h)$, then $h(x_{0})\neq h(x_{k})$. Clearly, ${\mathcal{H}}_{i}$ is closed under restrictions and non-empty. Now suppose that $h\in{\mathcal{H}}_{i}$ is such that there exists a tuple $a$ in a relation of $\mathfrak{A}$ whose entries contain $\operatorname{dom}(h)$. Let $S\subseteq A$ be a superset of $\operatorname{dom}(h)$ of cardinality at most $k$. If this relation is $D$, then $a$ equals $(x,y)$ and we may clearly extend $h$ to some $h^{\prime}\in{\mathcal{H}}_{i}$ with $h^{\prime}(x)\neq h^{\prime}(y)$. Otherwise, the relation must be $R$ and $a$ equals $(x_{0},x_{k},y,t)$ for some variables $x_{0},x_{k},y,t$ from a subformula $\phi$ of the form $\phi_{j}(x,y)$ for some $j\leq i$. Note that $h(x_{0})=h(x_{k})=h(y)$ if $j<i$, and $h(x_{0})\neq h(x_{k})$ and $h(x_{k})\neq h(y)$ otherwise. It is easy to see that in both cases we may extend $h$ to some $h^{\prime}\in{\mathcal{H}}_{i}$ with domain $S$. Finally, if $h\in{\mathcal{H}}_{i}$ for $i\in\{1,\dots,q\}$ and $S\subseteq A$ with $|S|\leq k$ is such that $\operatorname{dom}(h)\subseteq S$, then ${\mathcal{H}}_{i-1}$ contains an extension $h^{\prime}$ of $h$ with domain $S$: if $x\in S$ appears in a copy of $\phi_{i-1}$ whose variables intersect $\operatorname{dom}(h)$, we set $h^{\prime}(x)$ to the value taken by these variables under $h$. If $\operatorname{dom}(h)$ contains all the variables $x_{r},\dots,x_{s},x_{s+1}$ from $\phi_{i}$ for some $0\leq r<s<k$ and $x$ appears in the subformula $\phi_{i-1}(x_{s},x_{s+1})$ and some variable of $\phi_{i-1}(x_{r},x_{r+1})$ appears in $\operatorname{dom}(h)$, then we set $h^{\prime}(x)$ to the value taken by this variable under $h$. Otherwise, pick $h^{\prime}(x)$ to be a new value. Hence, ${\mathcal{H}}_{0},{\mathcal{H}}_{1},\dots,{\mathcal{H}}_{q}$ is indeed a winning strategy for Duplicator for the $q$ nested width $m$ guarded $k$ pebble game on $(\mathfrak{A},\mathfrak{C})$.
@@
We already know that the complement of $\operatorname{CSP}(\mathfrak{C})$ is in Datalog (Corollary 6.43), but not in GQ+ (Proposition 6.49). Proposition 6.47 shows that $\operatorname{CSP}(\mathfrak{B})$ is in MSO; then, as MSO is closed under negation, the complement of $\operatorname{CSP}(\mathfrak{B})$ is MSO-expressible as well.
@@ 7 Conclusion, Open Problems, and Prospect
We provided a game-theoretic characterisation of the problems in Guarded Second-order Logic that are equivalent to a Datalog program. We also proved the existence of canonical Datalog programs for GSO sentences whose models are closed under homomorphisms. To prove these results, we showed that every class of finite $\tau$-structures in GSO whose complement is closed under homomorphisms is a finite union of CSPs.
Our results imply that the so-called universal-algebraic approach, which has eventually led to the classification of finite-domain CSPs in Datalog [4], can be applied to study problems that are simultaneously in GSO and in Datalog (see [39]). They might also pave the way towards a syntactic characterisation of GSO $\cap$ Datalog and of MSO $\cap$ Datalog; however, we showed that the class of nested guarded queries from [18] is not expressive enough for this purpose. See Figure 4 for an overview of the considered logic, their inclusions, and references to examples that show that the inclusions are strict.
SO
Inklusions
and separations.
Datalog GSO GQ+ MSO GQ Nemodeq Guarded Datalog Modeq Monadic Datalog
Expl. 3.20CSP$(K_{3})$Thm. 6.18Expl. 3.20Prop. 6.21Expl. 2.5Expl. 2.5Expl. 2.5Expl. 2.5Prop. 6.30Prop. 6.6CSP$(K_{3})$[42]Prop. 6.30Prop. 6.6
Figure 4: A Hasse Diagram of the logical formalisms considered in this article, ordered by expressive power. The red labels give references to statements or examples in this text that show that the respective inclusion is strict.
@@
    Is every CSP in MSO $\cap$ Datalog the CSP of a reduct of a finitely bounded homogeneous structure?

### v3

@@ 5.7 Separation in MSO
Since $\operatorname{dom}(h)\leq k$ it is impossible that $r=0$ and s=k1s=k-1; hence, if $x_{0}$ and $x_{k}$ are from $\operatorname{dom}(h)$, then $h(x_{0})\neq h(x_{k})$. Clearly, ${\mathcal{H}}_{i}$ is closed under restrictions and non-empty. Now suppose that $h\in{\mathcal{H}}_{i}$ is such that there exists a tuple $a$ in a relation of $\mathfrak{A}$ whose entries contain $\operatorname{dom}(h)$. Let $S\subseteq A$ be a superset of $\operatorname{dom}(h)$ of cardinality at most $k$. If this relation is $D$, then $a$ equals $(x,y)$ and we may clearly extend $h$ to some $h^{\prime}\in{\mathcal{H}}_{i}$ with $h^{\prime}(x)\neq h^{\prime}(y)$. Otherwise, the relation must be $R$ and $a$ equals $(x_{0},x_{k},y,t)$ for some variables $x_{0},x_{k},y,t$ from a subformula $\phi$ of the form $\phi_{j}(x,y)$ for some $j\leq i$. Note that $h(x_{0})=h(x_{k})=h(y)$ if $j<i$, and $h(x_{0})\neq h(x_{k})$ and $h(x_{k})\neq h(y)$ otherwise. It is easy to see that in both cases we may extend $h$ to some $h^{\prime}\in{\mathcal{H}}_{i}$ with domain $S$. Finally, if $h\in{\mathcal{H}}_{i}$ for $i\in\{1,\dots,q\}$ and $S\subseteq A$ with $|S|\leq k$ is such that $\operatorname{dom}(h)\subseteq S$, then ${\mathcal{H}}_{i-1}$ contains an extension $h^{\prime}$ of $h$ with domain $S$: if $x\in S$ appears in a copy of $\phi_{i-1}$ whose variables intersect $\operatorname{dom}(h)$, we set $h^{\prime}(x)$ to the value taken by these variables under $h$. If $\operatorname{dom}(h)$ contains all the variables $x_{r},\dots,x_{s},x_{s+1}$ from $\phi_{i}$ for some $0\leq r<s<k$ and $x$ appears in the subformula $\phi_{i-1}(x_{s},x_{s+1})$ and some variable of $\phi_{i-1}(x_{r},x_{r+1})$ appears in $\operatorname{dom}(h)$, then we set $h^{\prime}(x)$ to the value taken by this variable under $h$. Otherwise, pick $h^{\prime}(x)$ to be a new value. Hence, ${\mathcal{H}}_{0},{\mathcal{H}}_{1},\dots,{\mathcal{H}}_{q}$ is indeed a winning strategy for Duplicator for the $q$ nested width $m$ guarded $k$ pebble game on $(\mathfrak{A},\mathfrak{C})$.
Note that the following is a strengthening of Corollary 5.39.
@@
We already know that the complement of $\operatorname{CSP}(\mathfrak{C})$ is in Datalog (Proposition 5.41), but not in GQ+ (Proposition 5.47). Proposition 5.45 shows that $\operatorname{CSP}(\mathfrak{B})$ is in MSO; then, as MSO is closed under negation, the complement of $\operatorname{CSP}(\mathfrak{B})$ is MSO-expressible as well.
6 A coNP-complete CSP in MSO
How tame is the intersection of MSO and Datalog? We have seen that every CSP which is in MSO and in Datalog can be expressed as the CSP of an $\omega$-categorical structure (Corollary 3.10). It would be interesting to know whether this can be strengthened further by showing that these problems can even be expressed as CSPs for a particularly well-behaved class of structures, namely for reducts of finitely bounded homogeneous structures (the definition of this class can be found below). It has been conjectured that every CSP for a structure from this class is in P or NP-complete [20]; we will refer to this conjecture as the infinite-domain dichotomy conjecture. This conjecture generalises the finite-domain complexity dichotomy that was conjectured by Feder and Vardi [32] and proved by Bulatov [22] and by Zhuk [46].
Definition 6.1.
A relational structure is called homogeneous if every isomorphism between finite substructures can be extended to an automorphism.
Homogeneous structures with a finite relational signature (and all of their reducts) are $\omega$-categorical. Reducts $\mathfrak{B}$ of homogeneous structures with finite relational signature are a particularly well-behaved subclass of the class of $\omega$-categorical structures. See [4] for an example of a result that only holds for such $\mathfrak{B}$, but not for $\omega$-categorical structures in general. However, there are uncountably many countable homogeneous digraphs with pairwise distinct CSP, and it follows that there are homogeneous digraphs with undecidable CSPs [19].
To define a class of structures whose CSPs have better computational properties we need the following concept. The age of a structure $\mathfrak{B}$ is the class of all finite structures that embed into $\mathfrak{B}$. A structure $\mathfrak{B}$ is called finitely bounded if there exists a finite set $\mathcal{F}$ of finite structures such that a finite structure $\mathfrak{A}$ belongs to the age of $\mathfrak{B}$ if and only if no structure in $\mathcal{F}$} embeds into $\mathfrak{A}$.
Example 6.2.
The structure $\mathfrak{B}$ from Section 5.6 is finitely bounded and homogeneous. $\triangle$
Clearly, the CSP of finitely bounded structures $\mathfrak{B}$ is in NP. It is also easy to see that every CSP of a reduct of a finitely bounded structure is in NP as well.
Question 6.3.
Is it true that every CSP in MSO is the CSP for a reduct of a finitely bounded homogeneous structure?
In this section, we do not answer this question, but at least we can show that the class of CSPs in MSO is (under complexity-theoretic assumptions) larger than the class of CSPs for reducts of finitely bounded structures (see Section 3.1). In particular, the class of CSPs that can be expressed in MSO does not fall into the scope of the mentioned infinite-domain CSP dichotomy conjecture.
A tournament is a directed graph such that for any pair of vertices $u$ and $v$, there is an edge from $u$ to $v$ or an edge from $v$ to $u$, but not both. Let $\mathcal{T}=\{\mathfrak{T}_{2},\mathfrak{T}_{3},\dots\}$ be the set of Henson tournaments: the tournament $\mathfrak{T}_{n}$, for $n\geq 2$, has vertices $0,1,\dots,n+1$ and the following edges:
  • $(i,i+1)$ for $i\in\{0,\dots,n\}$;
  • $(0,n+1)$;
  • $(j,i)$ for $i+1<j$ and $(i,j)\neq(0,n+1)$.
Proposition 6.4.
$\operatorname{CSP}(\mathfrak{H})$ can be expressed in MSO.
Proof 6.5.
We have to find an MSO sentence $\Phi$ that holds on a given digraph $(V;E)$ if and only if $(V;E)$ does not embed any of the tournaments from $\mathcal{T}$. We specify an MSO $\{X,E\}$-sentence $\Phi$, for a unary relation symbol $X$, that is true on a finite $\{X,E\}$-structure $\mathfrak{S}$ if and only if $(X^{\mathfrak{S}};E^{\mathfrak{S}})$ is isomorphic to $\mathfrak{T}_{n}$, for some $n\geq 2$. In $\Phi$, we existentially quantify over
  • two vertices $s,t\in X$ (that stand for the vertex 0 and the vertex $n+1$ in $\mathfrak{T}_{n}$).
The formula $\Phi$ has the following conjuncts:
  1. 1.
    a first-order formula that states that $E$ defines a tournament on $X$;
  2. 2.
  3. 3.
  4. 4.
    $E(s,t)$, $E(s,a)$, $E(a,b)$, and $E(x,s)$ for all $x\in X\setminus\{a,t\}$;
  5. 5.
  6. 6.
We claim that the MSO sentence $\forall x\colon\neg E(x,x)\wedge\forall X\colon\neg\Phi$ holds on a finite digraph if and only if the digraph is loopless and does not embed $\mathfrak{T}_{n}$, for all $n\geq 3$. The forwards implication easily follows from the observation that if $(X;T)$ is isomorphic to $\mathfrak{T}_{n}$, for some $n\geq 2$, then $\phi$ holds; this is straightforward from the construction of $\Phi$: we set $s$ to 0, $t$ to $n+1$, we set $A$ to the set of even and $B$ to the set of odd numbers in $\{1,\dots,n+1\}$. Conversely, suppose that $\Phi$ holds. Then $(X;T)$ is a tournament. We construct an isomorphism $f$ from $(X;T)$ to $\mathfrak{T}_{|X|-1}$ as follows. Define $f(s)\coloneqq 0$, $f(a)\coloneqq 1$, and $f(b)=2$. Since $E(a,b)$, by item 5 there exists exactly one $a^{\prime}\in A$ such that $E(b,a^{\prime})$ and $E(a^{\prime},a)$. Define $f(a^{\prime})\coloneqq 3$. If $a^{\prime}=t$ then we have found an isomorphism with $\mathfrak{T}_{2}$. Otherwise, the partial map $f$ defined so far is an embedding into $\mathfrak{T}_{n}$ for some $n\geq 3$. Item 6 and $E(b,a^{\prime})$ imply that there exists exactly one $b^{\prime}\in B$ such that $E(a^{\prime},b^{\prime})$ and $E(b^{\prime},b)$, and we define $f(b^{\prime})\coloneqq 4$. Continuing in this manner, we eventually define $f$ on all of $X$ and find an isomorphism with $\mathfrak{T}_{|X|-1}$.
This shows that, unless NP = coNP, $\operatorname{CSP}(\mathfrak{H})$ cannot be expressed as $\operatorname{CSP}(\mathfrak{B})$ for some reduct of a finitely bounded structure, because such CSPs are in NP.
Corollary 6.6.
If NP $\neq$ coNP, then there are CSPs in MSO that cannot be expressed as CSPs of finitely bounded structures.
We do not know how to show this statement without complexity-theoretic assumptions, even if we just want to rule out that $\operatorname{CSP}(\mathfrak{H})$ can be expressed as $\operatorname{CSP}(\mathfrak{B})$ for some reduct of a finitely bounded homogeneous structure. However, using the tools from Section 4.2 we can rule out that the complement of $\operatorname{CSP}(\mathfrak{H})$ can be expressed in Datalog. Note that this insight is obtained without resorting to any complexity-theoretic assumptions.
Theorem 6.7.
Let $\Phi$ be the MSO sentence from the proof of Proposition 6.4. Then $\llbracket\Phi\rrbracket$ cannot be expressed in Datalog.
Proof 6.8.
We use Theorem 4.12; it suffices to find for all $\ell,k$ a finite model $\mathfrak{A}$ of $\Phi$ such that Duplicator wins the existential $(\ell,k)$-pebble game for $\llbracket\Phi\rrbracket$ on $\mathfrak{A}$. Simply choose $\mathfrak{A}$ to be the tournament $\mathfrak{T}_{k+1}$; then clearly $\mathfrak{A}\models\neg\Phi$. Note that $\operatorname{CSP}(\mathfrak{H})\cap\llbracket\Phi\rrbracket=\emptyset$, so it suffices to find a winning strategy $\mathcal{H}$} for Duplicator for the existential $(\ell,k)$-pebble game on $(\mathfrak{A},\mathfrak{H})$.
Let $\mathcal{H}$ be the set of all partial embeddings $h$ from $\mathfrak{A}$ to $\mathfrak{B}$ such that $\operatorname{dom}(h)\leq k$. Then $\mathcal{H}$ is clearly non-empty and closed under restrictions. To verify that $\mathcal{H}$ is a winning strategy for Duplicator we still need to show that for all $h\in{\mathcal{H}}$ with $\operatorname{dom}(h)=\{a_{k-d+1},\dots,a_{k}\}$ and for all $a_{1},\dots,a_{k-d}\in A$ there is an extension $h^{\prime}\in{\mathcal{H}}$ of $h$ such that $h^{\prime}$ is also defined on $a_{1},\dots,a_{k-d}$. Let $\mathfrak{A}^{\prime}$ be the substructure of $\mathfrak{A}$ with the domain $\{a_{1},\dots,a_{k}\}$. Since $|A^{\prime}|<|A|$, there exists an embedding of $\mathfrak{A}^{\prime}$ into $\mathfrak{H}$; we identify the elements of $A^{\prime}$ with their image under this embedding, so that $\mathfrak{A}^{\prime}$ is now a substructure of $\mathfrak{H}$, and $h$ is an isomorphism between finite substructures of $\mathfrak{H}$. By the homogeneity of $\mathfrak{H}$, there exists an automorphism $\alpha$ of $\mathfrak{H}$ which extends $h$. Then the restriction of $\alpha$ to $a_{1},\dots,a_{k}$ is the desired extension $h^{\prime}$ of $h$ in $\mathcal{H}$}.
@@ 7 Conclusion, Open Problems, and Prospect
In this article, we provided a game-theoretic characterisation of the problems in Guarded Second-order Logic that are equivalent to a Datalog program. We also proved the existence of canonical Datalog programs for GSO sentences whose models are closed under homomorphisms. To prove these results, we showed that every class of finite $\tau$-structures in GSO whose complement is closed under homomorphisms is a finite union of CSPs.
Our results imply that the so-called universal-algebraic approach, which has eventually led to the classification of finite-domain CSPs in Datalog [5], can be applied to study problems that are simultaneously in GSO and in Datalog (see [41]). They might also pave the way towards a syntactic characterisation of GSO $\cap$ Datalog and of MSO $\cap$ Datalog; however, we showed that the class of nested guarded queries from [21] is not expressive enough for this purpose. See Figure 4 for an overview of the considered logics, their inclusions, and references to examples that show that the inclusions are strict.
[…]
Figure 4: A Hasse Diagram of the logical formalisms considered in this article, ordered by expressive power. The red labels give references to propositions or examples in this text that show that the respective inclusion is strict.
@@
    Question 6.3: Is every CSP in MSO $\cap$ Datalog the CSP of a reduct of a finitely bounded homogeneous structure?